I have a problem that starts with this equation:
x^2-2x-3=0
The next part of the equation is to move the constant to the right side of the equation.
x^2-2x-3+3=0+3
x^2-2x=3
x^2-2(x)=3
x^2-2+2x=3+2
x=5

Answer by fcabanski(1391)

Check my profile and contact me for one on one online tutoring.


It looks like you were completing the square, but incorrectly.


Step 1 is divide by the coefficient of x^2. It's 1, so that doesn't change the equation.


Step 2 is move the constant to the other side. You did that correctly.


x^2 -2x = 3


Step 3 is take 1/2 the coefficient of x, square it and add that to both sides. The coefficient of x is -2, 1/2 that is -1, squared it's 1.


x^2 -2x + 1 = 4


Step 4 is factor the left side. It will always be x then the sign of the original coefficient of z which in this case is -, then the number you added to both sides in step 3 (in this example it's 1). That whole thing is squared.


(x-1)^2 = 4 Take the square root of both sides


x-1 = + or - sqrt(4) = + or - 2.

Add 1 to both sides.


x = 1 + or - 2 = 3 or -1.